OpenStudy (anonymous):

What is the simplified form of the quantity 4 x squared minus 20x plus 25 over the quantity 2x minus 5

OpenStudy (anonymous):

Can you put it in equations or draw?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\frac{ 4x^2-20x+25 }{ 2x-5 } \]

OpenStudy (anonymous):

\[\frac{ (4x^2-10x-10x+25) }{ 2x-5 } \frac{ 2x(x-5)-5(x-5) }{ }\]

OpenStudy (anonymous):

Would the final answer be \[2x-5\] with a restriction of x\[\neq-\frac{ 5 }{ 2 }\]?

OpenStudy (anonymous):

cancel out the 2x-5's 2x-5 is your final answer

OpenStudy (anonymous):

YUP!

OpenStudy (anonymous):

basically always factor everything you can then cancel out terms. 2x-5 is your final answer

OpenStudy (anonymous):

Got it @Athentsu :) ?

OpenStudy (anonymous):

Thanks guys! Yeah. Just to check though, restriction would be -5/2 or 5/2?

OpenStudy (anonymous):

2x-5=0 2x=5 x=5/2 is the only restriction ;)

OpenStudy (anonymous):

Great, i thought it was -5/2 :( thanks for checking it for me